OpenStudy (anonymous):

Suppose that for a function f(x) we know that f(3)=5 and f ' (3)=7 We also 'have that g(x)=x^2*f(x) Find the equation of the line tangent to g(x) at x=3

OpenStudy (anonymous):

\[g'(x)=2xf(x)+x^2f'(x)\] by the product rule so \[g'(3)=2\times 3\times f(3)+3^2f'(3)\] plug in the numbers and get out your slope

OpenStudy (anonymous):

My calculator keeps giving me error

OpenStudy (anonymous):

this is not a calculator exercise

OpenStudy (anonymous):

\[2\times 2\times5+9\times 7\] is the slope

OpenStudy (anonymous):

6f(3)+9f'(3)

OpenStudy (anonymous):

83

OpenStudy (anonymous):

i meant \[2\times 3\times 5+9\times 7\]

OpenStudy (anonymous):

93

OpenStudy (anonymous):

is the slope

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

the point is \((3,3^2\times 5)\)

OpenStudy (anonymous):

better known as \((3,45)\) last step is point - slope formula

OpenStudy (anonymous):

ok how do you do that

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

do you know how to find the equation of a line given a point and a slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if the slope is \(m\) and the point is \((x_1,y_1)\) then the equation for the line is \[y-y_1=m(x-x_1)\] this is called the point - slope formula

OpenStudy (anonymous):

what are the x and y values if you on;y have (3,45)

OpenStudy (anonymous):

yes, that is all you need, one point \((3,45)\) and the slope \(m=93\)

OpenStudy (anonymous):

ok so then would it be 45=93-3

OpenStudy (anonymous):

ok lets back up a second

OpenStudy (anonymous):

you are being asked for the equation of a line right?

OpenStudy (anonymous):

usually it would look like \(y=mx+b\)

OpenStudy (anonymous):

yes the tangent line at x=3

OpenStudy (anonymous):

so your answer needs to look like \(y=mx+b\)

OpenStudy (anonymous):

y=93x+42

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] is the point slope formula replace \(m\) by \(93\), \(x_1\) by \(3\) and \(y_1\) by \(45\)

OpenStudy (anonymous):

you get \[y-45=93(x-3)\]

OpenStudy (anonymous):

now you can solve this for \(y\) to make it look like \(y=mx+b\)

OpenStudy (anonymous):

so then are the y and x 0

OpenStudy (anonymous):

oh ok just a minute

OpenStudy (anonymous):

y=138(x-3)

OpenStudy (anonymous):

\[y = 93 x-234\]

OpenStudy (anonymous):

oh ok I added the wrong thing together

OpenStudy (anonymous):

multiply out using the distributive law, then add 45

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then the y=93x-234 is the equation correct

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw